this is just a test to see if you actually send an email back, because i have to research a website for my honors algebra 2 class, and i chose this site. If i give a good review, my whole class may start using this site, because i find it most helpful! my question is, what is the difference between rational and irrational numbers? Thanks
-Sean Mason

Answer by venugopalramana(3286)

whatever is your purpose,as long as it is to learn and genuine,we can help..
qa rational number is one which can be expressed as p/q...where p and q are integers..example...0.25=1/4...and 0.3333333....=1/3 etc..
an irrational number is one which cannot be expressed as above..examples...square root of 2,cuberoot of three etc..